2. Hearts plc currently uses an absorption costing model to allocate overheads to its three products X, Y and Z. It is currenPlease answer all parts
Answer #1

a) Cost per unit under absorption costing model

Total Annual Overhead Costs: $

Machine set up Costs   26,550

Machine running Costs 66,400

Procurement Costs 48,000

Delivery Costs 54,320

   195,270

Overhead absorption rate:

A B C Total

Production Volumes 15,000 12,000 18,000

Labour hours per unit 0.1 0.15 0.2

Total Labour hours 1,500    1,800 3,600 6,900

Therefore,overhead absorption rate =$195,270/6,900= $28.30 per hour

Cost per unit :

A B C
$ $ $
Raw materials ($1.20*2/3/4 kg) 2.40 3.60 4.80
Direct Labour ($14.80*0.1/0.15/0.2hrs) 1.48 2.22 2.96
Overhead ($28.30*0.1/0.15/0.2hrs) 2.83 4.25 5.66
Total Cost per unit 6.71 10.07 13.42

b) Cost per unit using ABC Costing

Cost Driver:

Cost pool Cost Driver

   Machine set up Costs   26,550 (36 production runs 16+12+8)

Machine running Costs 66,400 (32,100 machine hours 7,500+8,400+16,200)

Procurement Costs 48,000 (94 purchase orders 24+28+42)

Delivery Costs 54,320 (140 Deliveries 48+30+62)

   195,270

Cost per machine set up $26,550/36=$737.50

Cost per machine hours $66,400/32,100= $2.0685

Cost per order $48,000/94 =$510.6383

Cost per delivery $54,320/140 =$388

Allocation of overheads to each product:

A B C Total
$ $ $ $
Machine set up costs 11,800 8,850 5,900 26,550
Machine-running costs 15,514 17,375 33,510 66,400
Procurement costs 12,255 14,298 21,447 48,000
Delivery costs 18,624 11,640 24,056 54,320
Total 58,193 52,163 84,913 195,270
Number of Units produced 15,000 12,000 18,000
Overhead Cost per unit 3.88 4.35 4.72

Total Cost per unit:

A B C
$ $ $
Raw materials 2.40 3.60 4.80
Direct Labour 1.48 2.22 2.96
Overhead 3.88 4.35 4.72
Total Cost per unit 7.76 10.17 12.48

c) Using activity -based costing

The following things can be observed -

Product A

The cost per unit of Product A is under ABC is $7.76 and under absorption costing is $6.71 and is 16% higher under ABC. So the company should do something to either reduce its cost or increase the selling price . Machine set up cost is also high for product A in comparison to other products and this could be reduced. Delivery cost are also high and this could be reduced through more efficiency.

Product B

The difference between the two approach is just $0.07 and the selling price is also $12 so it does not whichever method of costing is applied and product B is profitable

Product C

The unit cost for C is 7% lower under ABC when compared to absorption costing. Under absorption costing it is incurring a loss of $0.42 per unit as $13 is the cost. The reason could be that the number of production runs and product tests carried on C is relatively low. This leads to lower apportionment of machine set up cost to C.

ABC is therefore very useful in identifying that product C is more profitable than A.